The intensity (I) of a radio signal varies inversely as the square of the distance (d) from the transmitter. If the intensity is 23 W/m^2 at a distance from the transmitter of 2 km, what is the intensity st a distance of 6 km?
Answer by mathmate(429)

Question:
The intensity (I) of a radio signal varies inversely as the square of the distance (d) from the transmitter. If the intensity is 23 W/m^2 at a distance from the transmitter of 2 km, what is the intensity st a distance of 6 km?

Solution:
Given:
I(d)=k/d^2 => k=I(d)*d^2
23=k/2^2=k/4
Need I(6).
k=I(6)*6^2=23*2^2
Solve for I(6)
I(6)=(23*2^2)/6^2=23/9
= 2.56 approx.

Answer: the intensity at 6km is 2.56 W/m^2